A machine can make two products : A and B. the time taken to make A is Ta(hours) and that to make B is Tb(hours),the profit of a unit product B is pb.write the objective function to maximise the total profit based on a 24 hour schedule and the constraints to optimise the production quantities of A and B.
Answer #1

ASSUME X1 IS THE PRODUCTION :QUANTITY OF PRODUCT A AND X2 IS QUANTITY OF PRODUCT B.

objective function:

MAXIMIZE PROFIT = paX1 + pbX2

SUBJECT TO CONSTRAINT -

X1Ta + X2Tb \leqslant 24

non negative

x1 \geq 0

x2 \geq 0